# ROSE: A Recognition-Oriented Speech Enhancement Framework in Air Traffic Control Using Multi-Objective Learning

### v1

@@ Abstract
Radio speech echo is a specific phenomenon in the air traffic control (ATC) domain, which degrades speech quality and further impacts automatic speech recognition (ASR) accuracy. In this work, a recognition-oriented speech enhancement (ROSE) framework is proposed to improve speech intelligibility and also advance ASR accuracy, which serves as a plug-and-play tool in ATC scenarios and does not require additional retraining of the ASR model. Specifically, an encoder-decoder-based U-Net framework is proposed to eliminate the radio speech echo based on the real-world collected corpus. By incorporating the SE-oriented and ASR-oriented loss, ROSE is implemented in a multi-objective manner by learning shared representations across the two optimization objectives. An attention-based skip-fusion (ABSF) mechanism is applied to skip connections to refine the features. A channel and sequence attention (CSAtt) block is innovatively designed to guide the model to focus on informative representations and suppress disturbing features. The experimental results show that the ROSE significantly outperforms other state-of-the-art methods for both the SE and ASR tasks. In addition, the proposed approach can contribute to the desired performance improvements on public datasets.
Introduction
In air traffic control (ATC), speech is a primary means to achieve air-ground communication between the air traffic controller (ATCO) and the aircrew. ATCO issues spoken ATC instructions via the very high frequency radio to provide required services for aircrew, whereas the aircrew reads back the received instructions promptly to ensure a consentient understanding of the flight operation. In this procedure, the automatic speech recognition (ASR) technique is a preferred solution to convert the spoken instruction into readable texts, based on which contextual ATC elements are extracted to obtain traffic dynamics and further support ATC operations (Oualil et al. 2017; Lin et al. 2021c).
Recently, enormous efforts have been made to achieve the ASR task (Badrinath and Balakrishnan 2022; Lin et al. 2021b, a). By analyzing the ASR results, volatile noise is a major issue to limit ASR performance in the ATC environment, referring to the environment noise, device noise, radio transmission noise, speech echo, etc. Unlike the additive non-stationary signal (i.e., environment, device, radio), the speech echo is a specific overlapping phenomenon dedicatedly generated by the ATC communication between the sent and received ATCO speech.
@@ Introduction
Due to the returned ATCO instructions, a temporal delay between the sent speech and the received speech exists in the combined speech, which is defined as speech echo. As shown in Figure 1 (b), the speeches (only from ATCO) collected from CWP are outlined in the spectrogram manners and the time offset ($\Delta$T) is presented in the received speech. Compared to uncombined signals (sent or received speech), the combined speech spectrogram is distorted by redundant echo interference of signal overlapping, which further degrades the perception and intelligibility of the ATCO speech. By investigating the ASR performance, the accuracy of the ATCO speech (with speech echo) is obviously lower than that of the aircrew speech (without speech echo). Therefore, compared to other additive non-stationary noises, due to the ATC specificities, it is believed that the ASR performance can be significantly advanced by tackling the speech echo to enhance the speech quality.
@@
Since the temporal delay is affected by many factors (distance between CWP and radio station, radio frequency, etc.), the value of $\Delta$T is varied between 10 and 200 milliseconds, which is distinct for each utterance. Therefore, the redundant effect of speech echo is considered as special coupling noise and is not the solo reverberation or echo even though it is called ”echo” in the ATC domain. In the rest of this paper, the ATCO speeches with echo are regarded as noisy speeches, while their corresponding echo-free counterparts are clean speeches.
Speech enhancement (SE) is applied to eliminate noise interference on speech signals to obtain clean speech with higher intelligibility, which has been successfully applied to many applications, such as hearing aids, voice conversion and ASR system. In the early period, traditional SE methods (Lim and Oppenheim 1978; Boll 1979) usually use heuristic or direct algorithms to achieve denoising tasks. With the development of deep learning techniques, deep neural networks (DNN) are introduced into SE research, which have shown potent noise reduction ability in complex situations.
DNN-based SE models are divided into time-frequency (T-F) and time domain methods. T-F domain methods convert the raw signal to corresponding spectral magnitude as input by using short-time Fourier transform (STFT), which have achieved desired performance on SE task (Hu et al. 2020; Zhang et al. 2020; Chen et al. 2022; Schröter et al. 2022). Although the magnitude represents T-F information of the signal, some limitations are existing in the T-F methods (Wang, He, and Zhu 2021). Both the magnitude and phase information or complex value require high complexity of the model, which remains some challenges.
Time domain methods directly enhance noisy speech from raw waveform because it contains the full information of the signal. Various works have been achieved in convolutional neural networks (CNN), recurrent neural networks and generative adversarial networks (GAN) (Pascual, Bonafonte, and Serra 2017; Chen and Wang 2017; Macartney and Weyde 2018; Defossez, Synnaeve, and Adi 2020; Wang, He, and Zhu 2021; Kim and Seo 2021; Park et al. 2022). The raw waveform has the most local structure (millisecond-level signal) with non-discretization properties, which contain acoustic features such as the speaker’s timbre and linguistic patterns that can be obtained in the form of phonemes (Rethage, Pons, and Serra 2018), inspiring us to use time domain method to achieve the denoising task in the ATC domain with the auditory and phonetic demands.
As the front-end processor of the ASR system, SE can provide enhanced signals to reduce the non-robustness of the ASR model to redundant noise. In general, the primary goal of the SE technique is to improve the perceptual quality and intelligibility of speech for human hearing by modifying the temporal and frequency distributions of speech features. Unfortunately, the perception-oriented SE methods inevitably suffer from an inferior ASR performance due to the different optimization targets between the SE and ASR task. The solution of retraining the ASR model requires extra computational resources and still faces great challenges for its generalizability on different speech corpora.
@@
To address this issue, in this work, a novel recognition-oriented speech enhancement (ROSE) framework is proposed to improve both speech intelligibility and ASR performance, which serves as a plug-and-play tool in ATC scenarios and does not require additional retraining of the ASR model. ROSE is implemented using an encoder-decoder-based U-Net architecture with skip connections in the time domain. To enhance the feature fusion, an attention-based skip-fusion (ABSF) mechanism is applied to mine shared features from encoders using an attention mask (Giri, Isik, and Krishnaswamy 2019), which prevents the model from concatenating the hierarchical features directly. To further enhance the feature spaces, a channel and sequence attention (CSAtt) network is innovatively designed to learn objective-oriented feature weights in dual parallel attention paths by referring to (Lan et al. 2020).
To achieve the recognition-oriented SE model, the multi-objective learning is applied to achieve joint optimization of SE and ASR objectives by introducing the objective-oriented loss function. The details are as follows:
  1. 1.
@@
    The SE is to generate similar T-F features for the noisy speeches by referring to their clean counterparts. The mean absolute error (MAE) loss in the time domain and log-scale STFT-magnitude loss (Arik, Jun, and Diamos 2019) in the frequency domain serve as the optimization targets for the SE task.
  2. 2.
@@
    The enhanced output of the SE model for noisy signals only obtains desired performance by generating similar ASR-related features to that of clean signals. To this end, the well-designed algorithms for handcrafted feature engineering, such as spectrogram, Mel-frequency cepstral coefficients (MFCC), etc., inspire us to construct the optimization targets to retain the ASR performance for the SE output based on Fourier-based features.
Figure 2: Overall network architecture of ROSE.
In summary, the main contributions are as follows:
  • An end-to-end deep architecture ROSE is proposed to achieve the SE task of the ATC domain, which addresses the radio speech echo and achieves the desired accuracy without retraining the existing ASR model.
  • Multi-objective learning is applied to learn the common representations across the two optimization objectives to achieve the recognition-oriented SE model by introducing the T-F loss of the SE task and Fourier-based loss of the ASR task.
@@
  • An ABSF block is applied to achieve feature fusion using an attention mask. A CSAtt block is designed to learn informative features and suppress disturbance factors.
  • The experiment results show that ROSE outperforms other state-of-the-art (SOTA) models on both real-world ATC corpus and the public open-source datasets for SE and ASR tasks.
ROSE
@@
As shown in Figure 2 (a), the proposed framework is implemented using an end-to-end deep architecture, stacking multiple encoder and decoder layers with U-Net skip connections, in which the index of the decoder layer is the same as that of the corresponding encoder layer. The bidirectional long short-term memory (LSTM) is designed to capture the temporal dependencies among the frames and mine the correlations of signals in the past and future dimensions.
@@
The basic encoder block is comprised of sequential CNN blocks ([strided 1d-convolution (Conv1d), rectified linear unit (ReLU), 1$\times$1 Conv1d, gated linear unit (GLU)]) and CSAtt blocks. To formulate a U-Net shape, the basic decoder block applies an architecture similar to the encoder block, but with a different stacking order, i.e. ([1$\times$1 Conv1d, GLU, transposed 1d-convolution (ConvTr1d), ReLU]).
Attention-based Skip-fusion
@@
In general, skip-connection is to transfer shallow encoder features into decoder modules to formulate robust features. However, current practices fuse the encoder and decoder features by addition and concatenation operations with the same weight, which may not be optimal for complex tasks. In this study, the ABSF mechanism is designed to optimize the learned features, in which the gate block is to extract involved correlations from encoded features by multiplying them with an attention mask.
@@
As shown in Figure 2 (b), let the learned feature of the $i$th encoder layer be $E_{i}$, and the learned feature of the $i+1$th decoder layer be $D_{i+1}$. As shown in Equation 1, the convolution operation is applied to the above two feature maps, and the concatenation operation is performed to initialize the fusion attention coefficient matrix $B_{i}$.
$B_{i}=\sigma\left(\psi_{C}^{[1\times 1]}\left(E_{i}\right)\oplus\psi_{C}^{[1\times 1]}\left(D_{i+1}\right)\right)$ (1)
@@
where $\psi_{C}^{[K\times S]}$ are convolution operations with kernel size $K$ and stride $S$ of $C$ filters, $\oplus$ denotes the elementwise adding operation and $\sigma$ denotes the sigmoid activation. Note that the output feature dimension of the fusion module is the same as that of the input.
Based on the $B_{i}$, an attention mask $A_{i}$ is generated by the convolution network and the selected activation. Finally, the output feature $X_{i}$ is generated by Equation 2, in which $\otimes$ denotes term-wise product operation.
[…]
The Channel and Sequence Attention

### v2

@@ Abstract
Radio speech echo is a specific phenomenon in the air traffic control (ATC) domain, which degrades speech quality and further impacts automatic speech recognition (ASR) accuracy. In this work, a time-domain recognition-oriented speech enhancement (ROSE) framework is proposed to improve speech intelligibility and also advance ASR accuracy based on convolutional encoder-decoder-based U-Net framework, which serves as a plug-and-play tool in ATC scenarios and does not require additional retraining of the ASR model. Specifically, 1) In the U-Net architecture, an attention-based skip-fusion (ABSF) module is applied to mine shared features from encoders using an attention mask, which enables the model to effectively fuse the hierarchical features. 2) A channel and sequence attention (CSAtt) module is innovatively designed to guide the model to focus on informative features in dual parallel attention paths, aiming to enhance the effective representations and suppress the interference noises. 3) Based on the handcrafted features, ASR-oriented optimization targets are designed to improve recognition performance in the ATC environment by learning robust feature representations. By incorporating both the SE-oriented and ASR-oriented losses, ROSE is implemented in a multi-objective learning manner by optimizing shared representations across the two task objectives. The experimental results show that the ROSE significantly outperforms other state-of-the-art methods for both the SE and ASR tasks, in which all the proposed improvements are confirmed by designed experiments. In addition, the proposed approach can contribute to the desired performance improvements on public datasets.
Index Terms:
Speech enhancement, automatic speech recognition, attention-based skip-fusion, channel and sequence attention, multi-objective learning.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
In air traffic control (ATC), speech is a primary means to achieve air-ground communication between the air traffic controller (ATCO) and the aircrew. ATCO issues spoken ATC instructions via the very high frequency radio to provide required services for aircrew, whereas the aircrew reads back the received instructions promptly to ensure a consentient understanding of the flight operation. In this procedure, the automatic speech recognition (ASR) technique is a preferred solution to convert the spoken instruction into readable texts, based on which contextual ATC elements are extracted to obtain traffic dynamics and further support ATC operations [1, 2].
Recently, enormous efforts have been made to achieve the ASR task [3, 4, 5]. By analyzing the ASR results, volatile noise is a major issue that limits ASR performance in the ATC environment, referring to the environment noise, device noise, radio transmission noise, speech echo, etc. Unlike the additive non-stationary signal (i.e., environment, device, radio), the speech echo is a specific overlapping phenomenon dedicatedly generated by the ATC communication between the sent and received ATCO speech [6].
@@ I Introduction
Due to the returned ATCO instructions, a temporal delay between the sent speech and the received speech exists in the combined speech, which is defined as speech echo. As shown in Figure 1 (b), the speech (only from ATCO) collected from CWP is outlined in the spectrogram manner and the time offset ($\Delta$T) is presented in the received speech. Compared to uncombined signals (sent or received speech), the combined speech spectrogram is distorted by redundant echo interference of signal overlapping, which further degrades the perception and intelligibility of the ATCO speech.
By investigating the ASR performance, the accuracy of the ATCO speech (with speech echo) is obviously lower than that of the aircrew speech (without speech echo). Therefore, compared to other additive non-stationary noises, due to the ATC specificities, it is believed that the ASR performance can be significantly advanced by tackling the speech echo to enhance the speech quality.
@@
Since the temporal delay is affected by many factors (distance between CWP and radio station, radio frequency, etc.), the value of $\Delta$T is varied between 10 and 200 milliseconds, which is distinct for each utterance. Note that both the sent and received speech that comprise each combined speech are issued by the same ATCO and have the same text. The term “echo” describes the operation of a radio station returning a received speech with the same text as the sent speech. Although the type of signal that leads to the generation of the speech echo cannot be precisely defined, it can be treated as special coupling noise and is not the solo reverberation or echo even though it is called “echo” in the ATC domain.
As the front-end module of the ASR system, speech enhancement (SE) is applied to eliminate noise interference on speech signals to obtain clean speech with higher intelligibility. The enhanced clean speeches, i.e., the SE output, will subsequently serve as the input of the ASR system. In the rest of this paper, the ATCO speeches with echo are regarded as noisy speeches, while their corresponding echo-free counterparts are clean speeches.
However, previous works [7, 8] found that the enhanced signal from the SE model might yield inferior recognition accuracy for the subsequent ASR task. Some vital latent ASR-related information in the original noisy signal is diminished by the SE procedure with the redundant noise due to the different optimization targets between the SE and ASR tasks.
The strategy of joint optimization is an effective means. Some studies [9, 10] proposed a joint training approach to optimize the SE and ASR models together using enhanced signal only. Recent work [11] further combined enhanced and noisy signal features by utilizing an interactive fusion network to compensate for the neglected information in the original signal.
Although these methods have achieved good performance on SE and ASR tasks, they inevitably involve parameter modification of the ASR module. In the ATC domain, existing ASR models are trained with thousands of hours of speech corpora and have been deployed in over 20 real-world industrial ATC environments [5]. Therefore, retraining and redeploying the ASR models requires considering many issues, including data collecting, data annotation, extra computational resources and deployment costs. Moreover, the acquisition of high-quality speech is necessary to further support auditory-related downstream tasks in ATC scenarios such as emotion recognition and pilot/ATCO role identification.
@@
To address this issue, in this work, a novel recognition-oriented speech enhancement (ROSE) framework is proposed to improve both speech intelligibility and ASR performance, which serves as a plug-and-play tool in ATC scenarios and does not require additional retraining of the ASR model. Inspired by previous works [12, 13], ROSE is implemented using an encoder-decoder-based U-Net architecture with skip connections in the time domain. To enhance the model capacity for a multi-dimensional and robust feature space, two attention modules are designed to optimize the extracted feature representations. To achieve the recognition-oriented SE model, ASR-oriented optimization targets are designed to improve recognition performance in the ATC environment by learning robust feature representations, in which the multi-objective learning mechanism is applied to achieve joint optimization of SE and ASR objectives.
@@
Specifically, the SE is to generate similar time and frequency domain features for the noisy speeches by referring to their clean counterparts. Based on the previous works [12, 13], the mean absolute error (MAE) loss in the time domain and log-scale short-time Fourier transform (STFT)-magnitude loss [14] in the frequency domain serve as the optimization targets for the SE task.
@@
In addition, for a certain ASR model, the enhanced output of the SE model for noisy signals only obtains the desired performance by generating similar ASR-related features to that of clean signals. To this end, the well-designed algorithms for handcrafted feature engineering, such as spectrogram, Mel-frequency cepstral coefficients (MFCC), etc., inspire us to construct the optimization targets to retain the ASR performance for the SE output based on Fourier-based features.
In the encoder-decoder-based U-Net architecture, skip connections are built to generate fused features between encoder layers and corresponding decoder layers. To enhance the feature fusion, an attention-based skip-fusion (ABSF) module is applied to mine shared features from encoders using an attention mask, which prevents the model from concatenating the hierarchical features directly.
To enhance the feature spaces, a channel and sequence attention (CSAtt) module is designed to learn objective-oriented feature weights in dual parallel attention paths. The channel attention path focuses on the feature dimension to excite the informative points and suppress disturbance features. The sequence attention path is expected to allocate dedicated weights to highlight salient frames and discard echo frames, which further enhances the feature discriminative ability.
To advance the proposed SE model into ATC practices, a real-world monaural speech corpus collected from ATC scenarios is used to validate the proposed approach considering the dedicated generation mechanism of the ATC communication. The experimental results demonstrate that the proposed framework outperforms other comparative baselines, obtaining considerable improvements on both subjective and objective evaluation metrics of SE task. Most importantly, the enhanced speeches of the proposed model can achieve comparable ASR performance to clean speeches without any model retraining, i.e., 3.50% character error rate (CER) on the selected ASR model. In summary, this work contributes to the SE study in the ATC domain in the following ways:
1) To the best of our knowledge, this is the first work that applies the SE algorithm to address the speech echo in the ATC environment. A novel end-to-end deep architecture ROSE is proposed to achieve the SE and ASR tasks in the time domain, which obtains high-quality spoken signal and achieves the desired ASR accuracy without retraining the existing models. The framework can also be directly applied in a real-world ATC environment.
@@
2) In the proposed model, two attention modules are designed to optimize the extracted feature representations. An ABSF module is applied to skip connections to achieve feature fusion using an attention mask. A CSAtt module is innovatively designed to learn informative features and suppress disturbance factors from channel and sequence dimensions.
3) To improve recognition performance in the ATC environment, ASR-oriented objectives are designed to learn robust ASR features by optimizing the spectral distance between noisy and clean speeches. Moreover, a multi-objective learning mechanism is formulated to optimize the ROSE model using both the SE-related and ASR-related loss functions.
4) Extensive experimental results demonstrate that the ROSE outperforms other baselines on both SE and ASR tasks, and all the proposed technical improvements make expected contributions to related tasks. Most importantly, the proposed ROSE can be generalized to public open-source datasets and obtain comparable performance with state-of-the-art (SOTA) models.
II Related Works
II-A The ATC-Related ASR Works
In the ATC domain, the ASR technique aims at converting spoken ATC instructions into structured ATC-related elements [15], which is regarded as a key procedure to enhance the efficiency and safety of traffic operation [16]. Regarding multilingual ASR in the ATC scenarios (i.e., Chinese and English), a unified framework was constructed based on the acoustic model, pronunciation model and language model[2, 5]. The feature representation was studied to enhance the ASR performance [17]. Semi-supervised training strategy was presented to address the small sample problem in the ATC domain [18]. The transfer learning was introduced to enhance the model applicability for different scenes [4]. The contextual information of the real-time traffic operation was fused to improve the accuracy of ATC elements [1, 19].
The ASR technique is widely applied to enhance ATC applications, such as monitoring flight safety [16], reducing the controller’s workload [20], the ATCO training system [21] and decision support [22]. According to [5], it is believed that enhancing speech quality is a feasible solution to improve ASR performance, which is the primary focus of this work.
II-B The SE Works
In the past few decades, SE has been successfully applied to hearing aids [23], audio and video calls [24], voice conversion [25], speaker identification, and ASR system [26, 27].
Traditional SE methods [28, 29] usually use heuristic or direct algorithms to achieve denoising tasks. With the development of deep learning techniques, deep neural networks (DNN) are introduced into SE research, which have shown potent noise reduction ability in complex situations.
DNN-based SE models are divided into time-frequency (T-F) and time-domain methods. T-F domain methods convert the raw signal to the corresponding spectral magnitude as input by using STFT processing, which have achieved the desired performance on SE task [30, 31, 32]. Although the magnitude represents T-F information of the signal, some limitations are existing in the T-F methods [33]. Existing study [34] has proven that phase information is also crucial for enhancing speech quality. Recent works consider complex features and combine them with spectral magnitude or phase information to further achieve better performance [35, 36, 37, 38].
Time-domain methods directly enhance noisy speech from raw waveform in an end-to-end manner. Various works have been achieved in convolutional neural networks (CNN), recurrent neural networks and generative adversarial networks (GAN) [39, 12, 33, 40, 13]. Since the raw waveform contains the full information of the signal, the paradigm of learning directly from the waveform avoids the need for intermediate transformations (STFT and inverse STFT) and fully leverages the potentially valuable information (e.g., phase). Therefore, we propose the time-domain-based SE method to achieve the denoising task in the ATC domain.
Figure 2: Overall network architecture of ROSE and details of the ABSF module and CSAtt module.
III ROSE
III-A Problem Formulation
In general, the supervised SE task can be regarded as denoising the noisy speech signal to obtain an enhanced signal that is close to the reference signal. In the ATC domain, given the noisy combined speech with the speech echo $x=\{x_{1},x_{2},...,x_{T}\}$ composed of the received speech $s=\{s_{1},s_{2},...,s_{T}\}$, the sent speech with $\Delta T$ temporal delay $s^{{}^{\prime}}=\{s^{{}^{\prime}}_{1},s^{{}^{\prime}}_{2},...,s^{{}^{\prime}}_{T}\}$ and additional intrusive noise $n=\{n_{1},n_{2},...,n_{T}\}$ such that $x=s+s^{{}^{\prime}}+n$, the goal of the end-to-end SE task is to obtain the denoised signal $\tilde{s}$ directly from the noisy signal $x$, as Equation 1:
$\tilde{s}=\{\tilde{s}_{1},\tilde{s}_{2},...,\tilde{s}_{T}\}=f(x)$ (1)
where $T$ is the number of signal samples and $f(\cdot)$ denotes the SE model.
III-B Overview
@@
As shown in Figure 2 (a), the proposed framework is implemented using an end-to-end deep architecture, in which the backbone consists of a multi-layer encoder-decoder structure with U-Net skip connections by referring to [12]. In this work, the index of the decoder layer is the same as that of the corresponding encoder layer for convenient descriptions. The bidirectional long short-term memory (LSTM) is designed to capture the temporal dependencies among the frames and mine the correlations of signals in the past and future dimensions.
@@
The basic encoder block is comprised of sequential CNN blocks (1-d convolution (Conv1d) with kernel size $K$ and stride $S$, rectified linear unit (ReLU), pointwise Conv1d with kernel size 1 and stride 1, gated linear unit (GLU)) and CSAtt modules. To formulate a U-Net shape, the basic decoder block applies an architecture similar to the encoder block, but with a different stacking order, i.e. CSAtt modules and (pointwise Conv1d, GLU, transposed 1-d convolution (ConvTr1d), ReLU). The ABSF modules are integrated into the skip connections to achieve recalibration of features that are transmitted from encoder layers.
III-C Attention-based Skip-fusion
@@
In general, skip-connection is to transfer shallow encoder features into decoder blocks to formulate robust features. However, current practices fuse the encoder and decoder features by addition or concatenation operations with the same weight, which may not be optimal for complex tasks [39, 12]. In this study, the ABSF module is designed to optimize the learned features, in which the gate block is to extract involved correlations from encoded features by multiplying them with an attention mask.
@@
As shown in Figure 2 (b), let the learned feature of the $i$-th encoder layer be $E_{i}\in\mathbb{R}^{C\times L}$, and the learned feature of the $(i+1)$-th decoder layer be $D_{i+1}\in\mathbb{R}^{C\times L}$. As shown in Equation 2, the 1-d convolution layers are applied to the above two feature maps to halve the number of channels, and the concatenation operation is performed to initialize the fusion attention coefficient matrix $B_{i}$.
$B_{i}=\sigma\left(\psi_{\frac{C}{2}}^{[1,1]}\left(E_{i}\right)\oplus\psi_{\frac{C}{2}}^{[1,1]}\left(D_{i+1}\right)\right),B_{i}\in\mathbb{R}^{\frac{C}{2}\times L}$ (2)
@@
where $\psi_{C}^{[K,S]}$ are convolution operations with kernel size $K$ and stride $S$ of $C$ filters, $\oplus$ denotes the element-wise adding operation and $\sigma$ denotes the sigmoid activation.
Based on the $B_{i}$, a 1-d convolution layer followed by a sigmoid activation is used to project the number of channels as the same as the input shape, further obtaining an attention mask $A_{i}$. Finally, the output $X_{i}$ is obtained using the following rules, in which $\otimes$ denotes element-wise product operation.
$A_{i}=\sigma\left(\psi_{C}^{[1,1]}\left(B_{i}\right)\right),A_{i}\in\mathbb{R}^{C\times L}$ (3)
$X_{i}=D_{i+1}\oplus\left(E_{i}\otimes A_{i}\right),X_{i}\in\mathbb{R}^{C\times L}$ (4)
III-D The Channel and Sequence Attention